OpenStudy (anonymous):

A constant force is applied to an object, causing the object to accelerate at 10 m/s^2. What will the acceleration be if The object’s mass is halved?

OpenStudy (fifciol):

F=10m F=am/2 10m=am/2 a=20 so acceleration will be twice as before halved

OpenStudy (anonymous):

I got the same answer but it is wrong unfortunately, the correct answer is 30m/s^2 but I don't know how the book come up with is answer?

OpenStudy (anonymous):

This part b from the question. I answered a,c,d correctly but b I got 20m/s^2. the back of the book say the b is 30 m/s^2

OpenStudy (fifciol):

If you decrease the mass by certain amount the acceleration must go up by the same amount so that the product of them two(force) remains the same. It is said that the force is constant, so I have no idea why the book answer is 30

OpenStudy (doc.brown):

Fifciol is correct. Your book is incorrect. \[F=ma\]\[F_{1}=m(10m/s^{2})\]\[F_{2}=\frac{ 1 }{ 2}ma\]the force is constant, so\[F_{1} = F_{2}\]\[m(10m/s^{2}) = \frac{1}{2}ma\]\[\frac{m(10m/s^{2})}{m} = \frac{\frac{1}{2}ma}{m}\]\[10m/s^{2}= \frac{a}{2}\]\[2(10m/s^{2}) = 2\frac{a}{2}\]\[a = 20m/s^{2}\]

OpenStudy (anonymous):

Now three of us against the book. The book could be wrong I guess.

OpenStudy (xishem):

Maybe I should join the party. I agree that: \[a=20\frac{m}{s^2}\]

OpenStudy (anonymous):

I agree with all of you, but I have to verify.